An ab initio calculation of BH_{2}^{−} rovibronic energies: a very small singlet-triplet splitting

J.-P. Gu, G. Hirsch, R. J. Buenker, and \underline{Per Jensen}, FB 9 - Theoretische Chemie, Bergische Universität - Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany; P. R. Bunker, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6.

For each of the X̃^{3}B_{1}, ã^{1}A_{1}, and b̃^{1}B_{1} electronic states of the BH_{2}^{−} molecular ion we have calculated over 50 ab initio points on the potential energy surfaces using the MRD-CI method for all eight electrons. Analytical functions were fitted through the points for each state, and the functions for the ã and b̃ states were such that the surfaces remain degenerate at all linear configurations. For the X̃ state the rotation-vibration energies were calculated from the surface using the MORBID^{a} Hamiltonian, and for the ã and b̃ states the rovibronic energies were calculated using the RENNER^{b} Hamiltonian. For the X̃ state we obtain r_{e} = 1.2034 Å, α_{e} = 129.5^{∘} and a barrier to linearity of 2226 cm^{−1}. The vibrational wavenumbers are determined to be ν_{1} = 2353 cm^{−1}, ν_{2} = 883 cm^{−1} and ν_{3} = 2498 cm^{−1}. For the ã-state we determine r_{e} = 1.2369 Å, α_{e} = 105.7^{∘}, a barrier height to linearity of 6071 cm^{−1} and the vibrational wavenumbers ν_{1} = 2070 cm^{−1}, ν_{2} = 986 cm^{−1}, and ν_{3} = 2059 cm^{−1}. For the b̃-state we determine r_{e} = 1.1936 Å, α_{e} = 135.9^{∘}, a barrier height to linearity of 1069 cm^{−1} and the vibrational wavenumbers ν_{1} = 2399 cm^{−1}, ν_{2} = 915 cm^{−1}, and ν_{3} = 2591 cm^{−1}. The most interesting result we obtain is that the singlet-triplet splitting T_{0} (ã^{1}A_{1}) is only 153 cm^{−1} for ^{11}BH_{2}^{−}. Hence singlet-triplet perturbations will significantly affect all levels of the ground state.

aP. Jensen, J. Mol. Spectrosc. 128, 478–501 (1988).

bP. Jensen, M. Brumm, W. P. Kraemer, and P. R. Bunker, J. Mol. Spectrosc. 171, 31–57 (1995); M. Kolbuszewski, P. R. Bunker, W. P. Kramer, G. Osmann, and P. Jensen, Mol. Phys., in press.